OpenStudy (anonymous):

solve for theta on the interval [0, 360] 2 tan^2 theta-5 tan theta=0

OpenStudy (mendicant_bias):

This?\[2\tan ^{2}(\theta)-5\tan(\Theta) = 0\]

OpenStudy (anonymous):

\[\tan \theta=x \]

OpenStudy (anonymous):

2x^2-5x=0 x(2x-5)=0 x=0, 5/2 put the value of x & get the result.